Consider the acceleration of a particle along a straight line with an initial position of 15 m and an initial velocity of −16 m/s. What is the velocity at 1 s? Answer in units of m/s. What is the position at 1 s? Answer in units of m.
1
Expert's answer
2018-09-07T16:19:08-0400
The velocity at 1 s is
v=v_0+at=-16+1(1)=-15 m/s.
The position at 1 s is
s=s_0+v_0 t+(at^2)/2=15-16(1)+(1(1)^2)/2=-0.5 m.